# Bochner’s technique in Einstein’s non-symmetric geometry

### v1

@@ 3.4 The $f$-curvature operator
To simplify the calculations, in the rest of the paper we assume that the following Codazzi type condition is true:
$(\nabla^{g}_{{f}X}K)_{Y}Z=(\nabla^{g}_{{f}Y}K)_{X}Z,\quad X,Y,Z\in\mathfrak{X}_{M}.$ (52)
Here, $(\nabla^{g}_{{f}X}\,K)_{Y}\,Z=\nabla^{g}_{{f}X}(K_{Y}Z)-K_{\nabla^{g}_{{f}X}Y}\,Z-K_{Y}(\nabla^{g}_{{f}X}\,Z)$.
@@
For an Einstein’s connection $\nabla$ with conditions (27) and (52), we have
@@ 1. $(i)$
    ${R}^{f}_{X,Y}\,Z=R^{g}_{\,{f}X,{f}Y}\,Z+[K_{X},K_{Y}]Z-K_{[X,Y]_{f}}Z$;
    $({R}^{f}_{X,Y}\,\omega)(Z)=-\omega({R}^{f}_{X,Y}\,Z-[K_{X},K_{Y}]Z-K_{[X,Y]_{f}}Z)$;
    hence, $g({R}^{f}_{X,Y}\,Z,W)=-g({R}^{f}_{X,Y}\,W,Z)$,
@@ 3. $(iii)$
    $({R}^{f}_{X,Y}\,S)(Z_{1},\ldots,Z_{k})=(R^{g}_{\,{f}X,{f}Y}\,S)(Z_{1},\ldots,Z_{k})$
    $+([K_{X},K_{Y}]$ ${-}K_{[X,Y]_{f}})S(Z_{1},\ldots,Z_{k})-\sum\nolimits_{\,i}S(Z_{1},\ldots([K_{X},K_{Y}]{-}K_{[X,Y]_{f}})Z_{i},\ldots,Z_{k}).$ (53)
@@ 5. $(v)$
(i) Since ${f}{[X,Y]}_{f}=[{f}X,{f}Y]$ by the definition of ${\mathfrak{D}}^{f}$ and condition (27), we have
${R}^{f}_{X,Y}\,Z$ $=R^{g}_{\,{f}X,{f}Y}Z+(\nabla^{g}_{{f}X}\,K)_{Y}\,Z-(\nabla^{g}_{{f}Y}\,K)_{X}Z+[K_{X},K_{Y}]Z-K_{[X,Y]_{f}}Z,$
$({R}^{f}_{X,Y}\,\omega)(Z)$ $=-\omega(R^{g}_{\,{f}X,{f}Y}\,Z)+\omega((\nabla^{g}_{{f}X}\,K)_{Y}Z-(\nabla^{g}_{{f}Y}\,K)_{X}Z+[K_{X},K_{Y}]Z-K_{[X,Y]_{f}}Z).$
From this and (52) the first claim follows. Since (16) is true, that is, $K_{X},K_{Y}]:TM\to TM$ is skew-symmetric, ${R}^{f}_{X,Y}$ is also skew-symmetric: $g({R}^{f}_{X,Y}\,Z,W)=-g({R}^{f}_{X,Y}\,W,Z)$.
(ii) We calculate using our assumption (27), that is, ${\mathfrak{D}}^{f}=0$:
${R}^{f}_{X,Y}\,\psi={f}X({f}Y(\psi))-{f}Y({f}X(\psi))-({f}[X,Y]_{f})\psi={\mathfrak{D}}^{f}(X,Y)\psi=0.$
@@
$g({R}^{f}_{X,Y}\,Z,W)=g(R^{g}_{\,{f}X,{f}Y}\,Z,W)+g([K_{X},K_{Y}]Z-K_{[X,Y]_{f}}Z,\,W).$
Similarly, $g({R}^{f}_{X,Y}W,Z)=g(R^{g}_{\,{f}X,{f}Y}W,Z)+g([K_{X},K_{Y}]W-K_{[X,Y]_{f}}W,Z)$. By this and (51), we get
@@
$=(R^{g}_{\,{f}X,{f}Y}\,g)(Z,W)-g([K_{X},K_{Y}]Z,W)-g([K_{X},K_{Y}]W,Z)-g((K+K^{*})_{[X,Y]_{f}}Z,W)$
$=(R^{g}_{\,{f}X,{f}Y}\,g)(Z,W)-g((K+K^{*})_{[X,Y]_{f}}Z,W).$
Using $R^{g}_{\,{f}X,{f}Y}\,g=0$ and (15), we obtain ${R}^{f}_{X,Y}\,g=0$.
@@
(iv) The equality follows from (49) and (i).
(v) Since ${R}^{f}_{X,Y}\,Z=-{R}^{f}_{Y,X}\,Z$, see (i), the first equality follows. For the second one, we use (15):
$g({R}^{f}_{X,Y}Z,\,Z)=0;$
thus, the claim follows from the equality $g({R}^{f}_{X,Y}(Z+W),\,Z+W)=0$. ∎
@@
For an Einstein’s connection $\nabla$ satisfying (15), (27) and (52), we have
${\rm Ric}^{f}(X,Y)={\rm Ric}^{0}(X,Y)-g(K_{X}E+\sum\nolimits_{\,i}K_{[X,e_{i}]_{f}}e_{i},\,Y)+g(K_{X},\,K_{Y}+T(Y,\,\cdot\,),$ (54)
@@
${\rm Ric}^{f}(X,Y)$ $=\sum\nolimits_{\,i}{R}^{f}(X,e_{i},e_{i},Y)={\rm Ric}^{0}(X,Y)+\sum\nolimits_{\,i}g([K_{X},K_{e_{i}}]e_{i}-K_{[X,e_{i}]_{f}}e_{i},\ Y)$
$={\rm Ric}^{0}(X,Y)-g(K_{X}E,\,Y)+g(K_{X},\,K_{Y}+T(Y,\,\cdot\,)-\sum\nolimits_{\,i}K_{[X,e_{i}]_{f}}e_{i},\ Y),$
where $E$ is determined in (17). From the above the claim follows. ∎
Remark 3.
@@
The curvature operator $R^{g}_{X,Y}$ can be seen as a self-adjoint linear operator ${\cal R}^{g}$ on the space $\Lambda^{2}(M)$ of bivectors, e.g., [1, 8]. Similarly, we consider $R^{f}_{X,Y}$ as a linear operator, or as a corresponding bilinear form on $\Lambda^{2}(M)$. For this, using our assumption (16) that the operator $K_{X},K_{Y}]:TM\to TM$ is skew-symmetric, define a linear operator $\cal K$} on $\Lambda^{2}(M)$ by
$g({\cal K}(X\wedge Y),Z\wedge W):=g([K_{X},K_{Y}]Z-K_{[X,Y]_{f}}Z,\ W).$
@@
Note that if ${\cal P}\,{\cal R}^{g}+{\cal K}^{*}\neq{\cal R}^{g}\,{\cal P}+{\cal K}$, then ${\cal R}^{f}$ on $\Lambda^{2}(M)$ is not self-adjoint:
$({\cal R}^{f})^{*}=({\cal R}^{g}\,{\cal P}+{\cal K})^{*}={\cal P}\,{\cal R}^{g}+{\cal K}^{*}\neq{\cal R}^{g}\,{\cal P}+{\cal K}={\cal R}^{f}.$
@@ 3.5 The Weitzenböck $f$-curvature operator
Here, we use an Einstein’s connection $\nabla$ to introduce the central concept of the Bochner technique: the Weitzenböck $f$-curvature operator on tensors. Namely, we generalize the Weitzenböck curvature operator (3), for the case of a manifold $(M,G=g+F)$ with an $f$-connection $\nabla^{f}$.
@@
Define the Weitzenböck $f$-curvature operator on $(0,k)$-tensors $S$ over $(M,G=g+F)$ by
@@
$\thinspace\Re^{f}(S)(X_{1},\ldots,X_{k})=-2\sum\nolimits_{\,i,j,a;b<a}R^{f}(e_{i},X_{a},e_{j},X_{b})\cdot S(\underbrace{X_{1},\ldots,e_{j}}_{b},\underbrace{\ldots,e_{i}}_{a-b},\ldots,X_{k})$
$+\,\sum\nolimits_{\,i,a}{\rm Ric}^{f}(e_{i},X_{a})\cdot S(\underbrace{X_{1},\ldots,e_{i}}_{a},\ldots,X_{k}).$ (56)
@@
Let an Einstein’s connection $\nabla$ on a manifold $(M,G=g+F)$ satisfy conditions (15), (27) and (52). Then
@@
where the operator $\mathfrak{K}$ acts on forms $\omega\in\Lambda^{k}(M)$ by
$(\mathfrak{K}\,\omega)(X_{1},\ldots,X_{k})=2\sum\nolimits_{\,i,j,b<a}\big(g([K_{e_{i}},K_{X_{a}}]e_{j}-K_{[e_{i},X_{a}]_{f}}e_{j},\ X_{b})\big)\,\omega(\underset{b}{\underbrace{X_{1},\ldots,e_{j}}},\underset{a-b}{\underbrace{\ldots,e_{i}}}\ldots,X_{k})$
${-}\sum\nolimits_{\,a=1}^{k}\sum\nolimits_{\,i}\big(g(K_{e_{i}}E{+}\sum\nolimits_{\,j}\!K_{[e_{i},e_{j}]_{f}}e_{j},\,X_{a}){-}g(K_{e_{i}},\,K_{X_{a}}{+}T(X_{a},\,\cdot)\big)\,\omega(\underbrace{X_{1},\ldots,e_{i}}_{a},\ldots,X_{k}),$ (58)
for $k\geq 2$, and $(\mathfrak{K}\,\omega)(X){=}\sum\nolimits_{\,i}g(K_{e_{i}}E{+}\sum\nolimits_{\,j}\!K_{[e_{i},e_{j}]_{f}}e_{j},\,X){-}g(K_{e_{i}},\,K_{X}{+}T(X,\,\cdot))\omega(e_{i})$ for $k=1$.
@@
${R}^{f}(e_{i},X_{a},e_{j},X_{b})$ $={R}^{0}(e_{i},X_{a},e_{j},X_{b})+g([K_{e_{i}},K_{X_{a}}]e_{j}-K_{[e_{i},X_{a}]_{f}}e_{j},X_{b}),$
${\rm Ric}^{f}(e_{i},X_{a})$ $={\rm Ric}^{\,{0}}(e_{i},X_{a})-g(K_{e_{i}}E{+}\sum\nolimits_{\,j}K_{[e_{i},e_{j}]_{f}}e_{j},\,X_{a}){+}g(K_{e_{i}},\,K_{X_{a}}{+}T(X_{a},\,\cdot\,).$ (59)
Substituting (3.5) in (3.5) and using linearity in the curvature and ($(iii)$) yields (57) and (5). ∎
The next theorem generalizes (2) and [10, Theorem 2] to the case of an Einstein’s connection.
@@
Let an Einstein’s connection $\nabla$ satisfy conditions (15), (27) and (52). Then the following Weitzenböck type decomposition formula is valid for any form $\omega\in\Lambda^{k}(M)$:
@@
$=-\sum\nolimits_{\,j,a}((\nabla^{f})^{2}_{X_{a+1},e_{j}}\,\omega)\big(\underset{a+1}{\underbrace{X_{1},\ldots e_{j}}},\ldots,X_{k}\big)$
and
$\breve{\delta}^{f}\,d^{f}\omega(X_{1},\ldots,X_{k})=\nabla^{*{f}}(d^{f}\omega)(X_{1},\ldots,X_{k})=-\sum\nolimits_{\,j}\nabla_{e_{j}}^{f}(d^{f}\omega)(e_{j},X_{1},\ldots,X_{k})$
$=-\sum\nolimits_{\,j}\nabla^{f}_{e_{j}}\,\nabla^{f}_{e_{j}}\omega(X_{1},\ldots,X_{k})+\sum\nolimits_{\,j}\sum\nolimits_{a=0}^{k-1}(-1)^{a}\nabla^{f}_{e_{j}}\,\nabla^{f}_{X_{a+1}}\omega(e_{j},X_{1},\ldots,X_{a},X_{a+2},\ldots,X_{k})$
@@
$\Delta^{f}_{H}\,\omega=(\nabla^{*{f}}\,\nabla^{f}\omega)(X_{1},\ldots,X_{k})+\sum\nolimits_{\,j,a}\big((\nabla^{f})^{2}_{e_{j},X_{a+1}}\,\omega-(\nabla^{f})^{2}_{X_{a+1},e_{j}}\,\omega\big)(\underset{a+1}{\underbrace{X_{1},\ldots e_{j}}},\ldots,X_{k}).$
Thus, we have (60).
@@
For every bivector $X\wedge Y\in\Lambda^{2}(M)$, we consider a map ${\cal R}^{f}(X\wedge Y):\mathfrak{X}_{M}\to\mathfrak{X}_{M}$, given by
@@
$=$ $\hskip-0.28453ptR^{g}({f}X,{f}Y,Z,W)+g([K_{X},K_{Y}]Z,W),$
see Remark 3. Since bivectors are generators of the vector space $\Lambda^{2}(M)$, we obtain in this way a map ${\cal R}^{f}(\xi):\mathfrak{X}_{M}\to\mathfrak{X}_{M}$, where $\xi\in\Lambda^{2}(M)$ (similarly to algebraic curvature operator ${\cal R}^{g}(\xi)$).
@@
For an Einstein’s connection $\nabla$ with conditions (15), (27) and (52), the map ${\cal R}^{f}(\xi)$, where $\xi\in\Lambda^{2}(M)$, is skew-symmetric: $g({\cal R}^{f}(\xi)W,\,Z)=-g({\cal R}^{f}(\xi)Z,\,W)$.
@@
$g({\cal R}^{f}(X\wedge Y)Z,\ W)=R^{g}({f}X,{f}Y,Z,W)+g([K_{X},K_{Y}]Z-K_{[X,Y]_{f}}Z,\ W)$
$=-R^{g}({f}X,{f}Y,W,Z)-g([K_{X},K_{Y}]W-K_{[X,Y]_{f}}W,\ Z)=-g({\cal R}^{f}(X\wedge Y)W,\ Z).$
@@
For an Einstein’s connection $\nabla$ with conditions (15), (27) and (52), we have
@@
$=$ $\hskip-0.28453pt-\sum\nolimits_{\,\alpha}\big(g({\cal R}^{g}(\xi_{\alpha}){f}X,{f}Y)+g({\cal K}(X\wedge Y),\xi_{\alpha})\big)\xi_{\alpha},$
that completes the proof. ∎
@@
For an Einstein’s connection $\nabla$ with conditions (15), (27) and (52), if $S$ is a $(0,k)$-tensor on $(M,G=g+F)$, then
@@
We follow similar arguments as in the proof of [8, Lemma 9.3.3]:
@@
$=-\sum\nolimits_{\,i,\alpha}(\xi_{\alpha}S)(X_{1},\ldots,{\cal R}^{f}(\xi_{\alpha})X_{i},\ldots,X_{k})=-\sum\nolimits_{\,\alpha}({\cal R}^{f}(\xi_{\alpha})(\xi_{\alpha}S))(X_{1},\ldots,X_{k}).$`
Thus, (62)1 follows. Since ${\cal R}^{g}:\Lambda^{2}(M)\to\Lambda^{2}(M)$ is self-adjoint, there is a local orthonormal base $\{\xi_{a}\}$ of $\Lambda^{2}(M)$ such that ${\cal R}^{g}(\xi_{a})=\lambda_{a}\,\xi_{a}$. Using this base, for any $(0,k)$-tensors $S_{1}$ and $S_{2}$, we get
@@
Let $\nabla$ be an Einstein’s connection on a manifold $(M,G=g+F)$ satisfying conditions (15), (27) and (52).
@@
b) If $g({\cal R}^{f}(S),S)\geq-\varepsilon\,\|S\|^{2}$ for any $(0,k)$-tensor $S$, where $\varepsilon>0$, then $g({\Re}^{\,{f}}(S),S)\geq-\varepsilon\,C\,\|S\|^{2}$,
where a constant $C$ depends only on the type of $S$.
@@
Using (63) and a local orthonormal base $\{\xi_{\alpha}\}$ of $\Lambda^{2}(M)$ such that ${\cal R}^{g}(\xi_{\alpha})=\lambda_{\alpha}\xi_{\alpha}$, we get
@@
Let conditions (15), (27), (36) and (52) hold for an Einstein’s connection $\nabla$ on a closed manifold $(M,G=g+F)$. If $g({\cal R}^{f}(\omega),\omega)\geq 0$ for an f{f}-harmonic form $\omega\in\Lambda^{k}(M)$, then $\nabla^{f}\omega=0$.

### v2

@@ 3.4 The $f$-curvature operator
To simplify the calculations, define the following tensor:
$\Theta_{X,Y}:=(\nabla^{g}_{{f}X}K)_{Y}-(\nabla^{g}_{{f}Y}K)_{X}+[K_{X},K_{Y}]-K_{[X,Y]_{f}},\quad X,Y\in\mathfrak{X}_{M}.$
Here, $(\nabla^{g}_{{f}X}\,K)_{Y}Z=\nabla^{g}_{{f}X}(K_{Y}Z)-K_{\nabla^{g}_{{f}X}Y}Z-K_{Y}(\nabla^{g}_{{f}X}Z)$. If K(X,Y,Z)=K(X,Z,Y)K(X,Y,Z)=-K(X,Z,Y), see (15), then $\Theta_{X,Y}$ is skew-symmetric: $g(\Theta_{X,Y}Z,W)=-g(\Theta_{X,Y}W,Z)$.
@@
For an Einstein’s connection, we get
@@ 1. $(i)$
    ${R}^{f}_{X,Y}Z=R^{g}_{\,{f}X,{f}Y}Z+\Theta_{X,Y}Z$;  $({R}^{f}_{X,Y}\,\omega)(Z)=-\omega({R}^{f}_{X,Y}Z+\Theta_{X,Y}Z)$ for $\omega\in\Lambda^{1}(M)$;
    if (15) holds, then $g({R}^{f}_{X,Y}\,Z,W)=-g({R}^{f}_{X,Y}\,W,Z)$;
@@ 3. $(iii)$
    $\thinspace({R}^{f}_{X,Y}\,S)(Z_{1},\ldots,Z_{k})=$ $(R^{g}_{\,{f}X,{f}Y}\,S)(Z_{1},\ldots,Z_{k})+\Theta_{X,Y}S(Z_{1},\ldots,Z_{k})$
    $-\sum\nolimits_{\,i}S(Z_{1},\ldots\Theta_{X,Y}Z_{i},\ldots,Z_{k}).$ (52)
@@ 5. $(v)$
(i) By definition of ${R}^{f}_{X,Y}$, we get the first two equalities. Since $\Theta_{X,Y}:TM\to TM$ is skew-symmetric, ${R}^{f}_{X,Y}$ is also skew-symmetric.
(ii) We calculate
${R}^{f}_{X,Y}\,\psi={f}X({f}Y(\psi))-{f}Y({f}X(\psi))-({f}[X,Y]_{f})\psi={\mathfrak{D}}^{f}(X,Y)\psi.$
@@
$g({R}^{f}_{X,Y}\,Z,W)=g(R^{g}_{\,{f}X,{f}Y}\,Z,W)+g(\Theta_{X,Y}Z,\,W).$
Similarly, $g({R}^{f}_{X,Y}W,Z)=g(R^{g}_{\,{f}X,{f}Y}W,Z)+g(\Theta_{X,Y}W,Z)$. By this, (51) and $\nabla^{g}_{{\mathfrak{D}}^{f}(X,Y)}\,g=0$, we get
@@
$=(R^{g}_{\,{f}X,{f}Y}\,g)(Z,W)-g(\Theta_{X,Y}Z,W)-g(\Theta_{X,Y}W,Z)$
$=(R^{g}_{\,{f}X,{f}Y}\,g)(Z,W).$
Using $R^{g}_{\,{f}X,{f}Y}\,g=0$ in the above, we obtain ${R}^{f}_{X,Y}\,g=0$.
@@
(iv) The equalities follow from (49) and (i).
(v) Since ${R}^{f}_{X,Y}\,Z=-{R}^{f}_{Y,X}\,Z$, see (i), the first equality follows. For the second one, we use (15): $g({R}^{f}_{X,Y}Z,\,Z)=0$; thus, the claim follows from $g({R}^{f}_{X,Y}(Z+W),\,Z+W)=0$. ∎
@@
For an Einstein’s connection satisfying (15), we get
${\rm Ric}^{f}(X,Y)={\rm Ric}^{0}(X,Y)+g\big(\sum\nolimits_{\,i}\Theta_{X,e_{i}}e_{i},\,Y\big),$ (53)
@@
${\rm Ric}^{f}(X,Y)$ $=\sum\nolimits_{\,i}{R}^{f}(X,e_{i},e_{i},Y)={\rm Ric}^{0}(X,Y)+\sum\nolimits_{\,i}g(\Theta_{X,e_{i}}e_{i},\,Y).$
From the above the claim follows. ∎
@@
The curvature operator $R^{g}_{X,Y}$ can be seen as a self-adjoint linear operator ${\cal R}^{g}$ on the space $\Lambda^{2}(M)$ of bivectors, e.g., [1, 8]. Similarly, we consider $R^{f}_{X,Y}$ as a linear operator, or as a corresponding bilinear form on $\Lambda^{2}(M)$. For this, using assumption (15), define a linear operator $\cal K$} on $\Lambda^{2}(M)$ by
$g({\cal K}(X\wedge Y),Z\wedge W):=g(\Theta_{X,Y}Z,\ W).$
@@ 3.5 The Weitzenböck $f$-curvature operator
Here, we use an Einstein’s connection $\nabla$ to introduce the central concept of the Bochner technique: the Weitzenböck $f$-curvature operator on tensors, generalizing the operator (3), for the case of a manifold $(M,G=g+F)$ with an $f$-connection $\nabla^{f}$.
@@
Define the Weitzenböck $f$-curvature operator on $(0,k)$-tensors $S$ on a manifold $(M,G=g+F)$ by
@@
$\Re^{f}(S)(X_{1},\ldots,X_{k})=$ $-2\sum\limits_{\,i,j,a;\,b<a}R^{f}(e_{i},X_{a},e_{j},X_{b})\cdot S(\underbrace{X_{1},\ldots,e_{j}}_{b},\underbrace{\ldots,e_{i}}_{a-b},\ldots,X_{k})$
$+\,\sum\nolimits_{\,a=1}^{k}\sum\nolimits_{\,i}{\rm Ric}^{f}(e_{i},X_{a})\cdot S(\underbrace{X_{1},\ldots,e_{i}}_{a},\ldots,X_{k}).$ (55)
@@
Let an Einstein’s connection $\nabla$ on a manifold $(M,G=g+F)$ satisfy (15). Then
@@
where the operator $\mathfrak{K}$ acts on forms $\omega\in\Lambda^{k}(M)$ for $k\geq 2$ by
$(\mathfrak{K}\,\omega)(X_{1},\ldots,X_{k})=$ $\ 2\sum\nolimits_{\,i,j,\,b<a}g(\Theta_{e_{i},X_{a}}e_{j},\ X_{b})\,\omega(\underset{b}{\underbrace{X_{1},\ldots,e_{j}}},\underset{a-b}{\underbrace{\ldots,e_{i}}}\ldots,X_{k})$
$+\sum\nolimits_{\,a=1}^{k}\sum\nolimits_{\,i}g\big(\sum\nolimits_{\,j}\Theta_{e_{i},\,e_{j}}e_{j},\,X_{a}\big)\,\omega(\underbrace{X_{1},\ldots,e_{i}}_{a},\ldots,X_{k}),$ (57)
and $(\mathfrak{K}\,\omega)(X)=\sum\nolimits_{\,i}g\big(\sum\nolimits_{\,j}\Theta_{e_{i},\,e_{j}}e_{j},\,X\big)\,\omega(e_{i})$ for $k=1$.
@@
${R}^{f}(e_{i},X_{a},e_{j},X_{b})={R}^{0}(e_{i},X_{a},e_{j},X_{b})+g(\Theta_{e_{i},X_{a}}e_{j},X_{b}),$
$\thinspace{\rm Ric}^{f}(e_{i},X_{a})={\rm Ric}^{\,{0}}(e_{i},X_{a})+g(\sum\nolimits_{\,j}\Theta_{e_{i},\,e_{j}}e_{j},\,X_{a}).$ (58)
Using (3.5) in (3.5), by linearity in the curvature and ($(iii)$) we get (56) and (6). ∎
We generalizes (2) and [10, Theorem 2] to the case of an Einstein’s connection.
@@
Let an Einstein’s connection $\nabla$ satisfy (15). Then the Weitzenböck type decomposition formula is valid for any $\omega\in\Lambda^{k}(M)$:
@@
$=-\sum\nolimits_{\,j,a}((\nabla^{f})^{2}_{X_{a+1},e_{j}}\,\omega)\big(\underset{a+1}{\underbrace{X_{1},\ldots e_{j}}},\ldots,X_{k}\big),$
$\breve{\delta}^{f}\,d^{f}\omega(X_{1},\ldots,X_{k})$ $=\nabla^{*{f}}(d^{f}\omega)(X_{1},\ldots,X_{k})$
$=-\sum\nolimits_{\,j}\nabla_{e_{j}}^{f}(d^{f}\omega)(e_{j},X_{1},\ldots,X_{k})=-\sum\nolimits_{\,j}\nabla^{f}_{e_{j}}\,\nabla^{f}_{e_{j}}\omega(X_{1},\ldots,X_{k})$
$+\sum\nolimits_{\,j}\sum\nolimits_{a=0}^{k-1}(-1)^{a}\nabla^{f}_{e_{j}}\,\nabla^{f}_{X_{a+1}}\omega(e_{j},X_{1},\ldots,X_{a},X_{a+2},\ldots,X_{k})$
@@
$\Delta^{f}_{H}\,\omega=(\nabla^{*{f}}\nabla^{f}\omega)(X_{1},\ldots,X_{k}){+}\sum\limits_{\,j,a}\big((\nabla^{f})^{2}_{e_{j},X_{a+1}}\,\omega-(\nabla^{f})^{2}_{X_{a+1},e_{j}}\,\omega\big)(\underset{a+1}{\underbrace{X_{1},\ldots e_{j}}},\ldots,X_{k}).$
Thus, we have (59). ∎
@@
For a bivector $X\wedge Y\in\Lambda^{2}(M)$, consider a map ${\cal R}^{f}(X\wedge Y):\mathfrak{X}_{M}\to\mathfrak{X}_{M}$, given by
@@
$=$ $\hskip-0.28453ptR^{g}({f}X,{f}Y,Z,W)+g(\Theta_{X,Y}Z,W),$
see (iv) of Proposition 7. Since bivectors are generators of the vector space $\Lambda^{2}(M)$, we get a map ${\cal R}^{f}(\xi):\mathfrak{X}_{M}\to\mathfrak{X}_{M}$, where $\xi\in\Lambda^{2}(M)$ (similarly to algebraic curvature operator ${\cal R}^{g}(\xi)$).
@@
For an Einstein’s connection $\nabla$ satisfying (15), the map ${\cal R}^{f}(\xi)$, where $\xi\in\Lambda^{2}(M)$, is skew-symmetric: $g({\cal R}^{f}(\xi)W,\,Z)=-g({\cal R}^{f}(\xi)Z,\,W)$.
@@
$g({\cal R}^{f}(X\wedge Y)Z,\ W)=R^{g}({f}X,{f}Y,Z,W)+g(\Theta_{X,Y}Z,\ W)$
$=-R^{g}({f}X,{f}Y,W,Z)-g(\Theta_{X,Y}W,\ Z)=-g({\cal R}^{f}(X\wedge Y)W,\ Z).$
@@
Let $\nabla$ be an Einstein’s connection satisfying (15). Then
@@
$=$ $\hskip-0.28453pt-\sum\nolimits_{\,\alpha}\big(g({\cal R}^{g}(\xi_{\alpha}){f}X,{f}Y)+g({\cal K}(X\wedge Y),\xi_{\alpha})\big)\xi_{\alpha},\hskip 85.35826pt\Box$
@@
Let $\nabla$ be an Einstein’s connection satisfying (15), and $S$ a $(0,k)$-tensor on $(M,G=g+F)$. Then
@@
We prove (61)1 using similar arguments as in the proof of [8, Lemma 9.3.3]:
@@
$=-\sum\nolimits_{\,i,\alpha}(\xi_{\alpha}S)(X_{1},\ldots,{\cal R}^{f}(\xi_{\alpha})X_{i},\ldots,X_{k})=-\sum\nolimits_{\,\alpha}({\cal R}^{f}(\xi_{\alpha})(\xi_{\alpha}S))(X_{1},\ldots,X_{k}).$
Since ${\cal R}^{g}:\Lambda^{2}(M)\to\Lambda^{2}(M)$ is self-adjoint, there is a local orthonormal base $\{\xi_{a}\}$ of $\Lambda^{2}(M)$ such that ${\cal R}^{g}(\xi_{a})=\lambda_{a}\,\xi_{a}$. For this base and any $(0,k)$-tensors $S_{1}$ and $S_{2}$, we get
@@
Let $\nabla$ be an Einstein’s connection on a manifold $(M,G=g+F)$ satisfying (15).
@@
b) If $g({\cal R}^{f}(S),S)\geq-\varepsilon\,\|S\|^{2}$ for any $(0,k)$-tensor $S$, where $\varepsilon>0$, then
$g({\Re}^{\,{f}}(S),S)\geq-\varepsilon\,C\,\|S\|^{2}$, where a constant $C$ depends only on the type of $S$.
@@
By (62), for a local orthonormal base $\{\xi_{\alpha}\}$ of $\Lambda^{2}(M)$ such that ${\cal R}^{g}(\xi_{\alpha})=\lambda_{\alpha}\xi_{\alpha}$, we get
@@
Let conditions (15) and (36) hold for an Einstein’s connection $\nabla$ on a closed manifold $(M,G=g+F)$. If $g({\cal R}^{f}(\omega),\omega)\geq 0$ for an f{f}-harmonic form $\omega\in\Lambda^{k}(M)$, then $\nabla^{f}\omega=0$.